the sum of four consecutive odd integers 216.find the four integers.
Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

the sum of four consecutive odd integers 216.find the four integers.
------------------------------------
1st 2x-3
2nd 2x-1
3rd 2x+1
4th 2x+3
------------------
EQUATION:
8x = 216
x = 27
------------
1st : 2x-3 = 54-3 = 51
2nd: 53
3rd: 55
4th: 57
==================
Cheers,
Stan H.

Answer by checkley77(12844)

X+X+2+X+4+X+6=216
4X+12=216
4X=216-12
4X=204
X=204/4
X=51 FOR THE LOWEST INTEGER.
51+2=53 FOR THE NEXT INTEGER.
51+4=55 FOR THE THIRD INTEGER
51+6=57 FOR THE LARGEST INTEGER.
PROOF:
51+53+55+57=216
216=216